RC Meeting and Runplan, Friday, August 24

From HallCWiki
Jump to: navigation, search
  • Status
    • LN2 supply issue addressed
    • Hall locking up today
    • Target cooled down, liquid hydrogen in loop 2
    • Beamline vacuum improving, good enough for beam delivery
    • First beam to hall potentially swing/owl
    • SHMS steering test will be done at 7 degrees instead of 6 - when we need to go to 6 degrees, we will need to redo the test

  • Run plan
    • Friday SWING-Saturday OWL and DAY
      • Initial beam setup to the hall
      • If MCC asks, the Compton electron detector is OUT
      • Make sure target is OUT
      • After initial setup - SHMS steering plan will be executed (please call RC if not already on site). When MCC asks, ramp up magnets (one at a time) to settings for -6.053 GeV.
      • After completion of SHMS steering plan, proceed with detector and trigger checkout